Sound with a frequency of 650Hz from a distant source passes through a doorway. 1.10m wide in a sound absorbing wall. Find

(a) the number and (b) the angular directions of the diffraction minima at listening positions along a line parallel to the wall.


1
Expert's answer
2021-03-01T07:21:34-0500

Answer

Wavelength

"\lambda=\frac{v_s}{f}=\frac{343}{650}=0.528m"

Conditions for diffraction minima

"asin\theta=m\lambda"

Where

"m=\pm1 ,\pm2 ,\pm3,....."

Now

"\theta_1=\sin^{-1}(\pm1\frac{0.528}{1.1}) =\pm28.7^°"


"\theta_2=\sin^{-1}(\pm2\frac{0.528}{1.1}) =\pm73.7^°"



"\theta_3=\sin^{-1}(\pm3\frac{0.528}{1.1}) =" non-exist


Therefore

a) four minima

b) angular directions are

"\pm28.7^°, \pm73.7^°."